OpenStudy (anonymous):

300 room hotel in Macau is filled to full capacity every night at $800 per room. For each additional $10 increase in rent, 3 fewer rooms are rented. If each rented room costs $160 to service per day, how much should the hotel charge for each room to maximize profit? Please help me, I got stuck in formulating these variables Thank You

ganeshie8 (ganeshie8):

lets label the variables : P : profit x : number of rooms filled w : rent of room

ganeshie8 (ganeshie8):

P = wx - 160w ---------------- (1)

ganeshie8 (ganeshie8):

x = 300 - (w-800)/10 --------------------(2)

ganeshie8 (ganeshie8):

i think now u can substitute x in (1). and maximize P

ganeshie8 (ganeshie8):

lol typo in equation (2) it must be : x = 300 - 3(w-800)/10 --------------------(2)

OpenStudy (anonymous):

Thank You so much for your help

ganeshie8 (ganeshie8):

welcome.. :))

OpenStudy (anonymous):

sorry to disturb, but in equation 2 didn;t you mean p=wx-160x, thank you

ganeshie8 (ganeshie8):

yep... another typo lol. cost = 160 times num.of.rooms nice catch :)

ganeshie8 (ganeshie8):

P = wx - 160x ---------------- (1) x = 300 - 3(w-800)/10 --------------------(2) now both look correct to me..

OpenStudy (anonymous):

thx mate

ganeshie8 (ganeshie8):

yw :)